# UniPruning: Unifying Local Metric and Global Feedback for Scalable Sparse LLMs

### v1

@@ 5.3 Ablation Study
Local Metric. To evaluate the sensitivity of pruning performance to the choice of local saliency metric, we conduct ablation studies comparing different saliency metric designs. Specifically, we experiment with magnitude-based scoring, activation-aware scoring as used in Wanda (Sun et al., 2024), and the combination of row/column norm-based relative importance with activation signals as in RIA (Zhang et al., 2024). Furthermore, we incorporate the stochastic variant of RIA (stochRIA) proposed by (Yi & Richtárik, 2025), which introduces randomness into the scoring process to mitigate biases introduced by deterministic saliency measures and to enhance exploration during pruning.
@@ A.4.2 Relative Error Property
By (LABEL:Eq:subgradbound-Pk) and Lemma 2(iv), $\frac{1}{K}\sum_{k=1}^{K}\|H_{k}\|^{2}\rightarrow 0$ at a rate of ${\cal O}(1/K)$.

### v2

@@ 5.4 Ablation Study
Local Metric. To evaluate the sensitivity of pruning performance to the choice of local saliency metric, we conduct ablation studies comparing different saliency metric designs. Specifically, we experiment with magnitude-based scoring, activation-aware scoring as used in Wanda (sun2024simple), and the combination of row/column norm-based relative importance with activation signals as in RIA (zhang2024plug). Furthermore, we incorporate the stochastic variant of RIA (stochRIA) proposed by (yi2025symmetric), which introduces randomness into the scoring process to mitigate biases introduced by deterministic saliency measures and to enhance exploration during pruning.
@@ A.4.2 Relative Error Property
By Lemma 2(iv), $\frac{1}{K}\sum_{k=1}^{K}\|H_{k}\|^{2}\rightarrow 0$ at a rate of ${\cal O}(1/K)$.